# CooPre: Cooperative Pretraining for V2X Cooperative Perception

### v1

@@ Abstract
Existing Vehicle-to-Everything (V2X) cooperative perception methods rely on accurate multi-agent 3D annotations. Nevertheless, it is time-consuming and expensive to collect and annotate real-world data, especially for V2X systems. In this paper, we present a self-supervised learning method for V2X cooperative perception, which utilizes the vast amount of unlabeled 3D V2X data to enhance the perception performance. Beyond simply extending the previous pre-training methods for point-cloud representation learning, we introduce a novel self-supervised Cooperative Pretraining framework (termed as CooPre) customized for a collaborative scenario. We point out that cooperative point-cloud sensing compensates for information loss among agents. This motivates us to design a novel proxy task for the 3D encoder to reconstruct LiDAR point clouds across different agents. Besides, we develop a V2X bird-eye-view (BEV) guided masking strategy which effectively allows the model to pay attention to 3D features across heterogeneous V2X agents (i.e., vehicles and infrastructure) in the BEV space. Noticeably, such a masking strategy effectively pretrains the 3D encoder and is compatible with mainstream cooperative perception backbones. Our approach, validated through extensive experiments on representative datasets (i.e., V2X-Real, V2V4Real, and OPV2V), leads to a performance boost across all V2X settings. Additionally, we demonstrate the framework’s improvements in cross-domain transferability, data efficiency, and robustness under challenging scenarios. The code will be made publicly available.
@@ I Introduction
Figure 1: Illustration of our CooPre framework. Given multi-agent LiDAR input, CooPre pretrains the 3D encoder via a LiDAR point cloud reconstruction task. During the finetuning stage, the initialized 3D encoder will be used for 3D detection task.
In this paper, based on the aforementioned observations, we propose a simple yet effective self-supervised multi-agent pretraining framework named CooPre, showing that Cooperative Pretraining enables the model to learn meaningful prior representation of the holistic 3D environment before the perception task, as illustrated in Fig.1. Our framework leverages the benefits of unlabeled LiDAR point cloud data transmitted from different agents, allowing the model to reconstruct the point cloud location and learn essential prior knowledge of scenarios (e.g., intersections or corridors) and LiDAR sensor distributions (e.g., range, placement position, and sparsity) of each agent from a bird-eye-view (BEV) perspective. This design is compatible with current cooperative perception backbones. It makes the reconstruction task more challenging due to the enlarged perception field, which helps mitigate issues related to sparse feature points in far-range and occlusion scenarios. Compared to the other two training paradigms in Fig. 2, our pretraining framework is annotation-free, which is the main contribution to representation learning in V2X cooperative perception.
Following the introduction of this straightforward approach, the remainder of the contribution of this paper focuses on explaining how, where, and why this approach works in V2X cooperative perception scenarios. Our analysis demonstrates: 1) why pretraining by showcasing our framework’s strong generalizability to unseen distributions, which is beneficial for domain adaptation, especially when the target dataset is limited in scale; 2) why reconstruction pretraining by highlighting our framework’s attention to the geometrical and topological features of rigid-body objects (e.g., cars and trucks); 3) why multi-agent pretraining by illustrating improvements in far-range perception and occlusion handling compared to train-from-scratch and single-agent pretraining frameworks[18, 19, 20]. Through extensive ablation studies and experiments on V2X-Real[1], V2V4Real[8], and OPV2V[21] datasets, we demonstrate the effectiveness of CooPre and shed light on how it works.
@@ II-A Cooperative Perception
Single-vehicle systems struggle with occlusions and long-distance perception in complex traffic environments due to the limitation of the perception range of LiDAR devices. Cooperative systems, on the other hand, enhance detection performance by sharing raw data (Early Fusion), detection outputs (Late Fusion), or intermediate bird-eye-view (BEV) representations (Intermediate Fusion) among connected agents [21, 22, 23]. With the recent advancement of a variety of simulation and real-world dataset curations[21, 8, 3, 24, 4, 1, 5], many literature[21, 2, 25, 22, 26, 23, 27] have been discussing the algorithmic designs of collaborative modes from vehicle-to-vehicle (V2V) collaboration to vehicle-to-everything (V2X) collaboration in different traffic scenarios. Noticeably, the intermediate fusion strategy has been the primary direction since it achieves the best trade-off between accuracy and bandwidth requirements. After applying a 3D encoder[28, 29, 30] to the input LiDAR feature, the intermediate fusion strategy involves projecting 3D features to BEV features where agents will perform interactions before final detection results. For example, AttFuse[21] utilizes a simple agent-wise single-head attention to fuse all features, whereas V2X-ViT[2] presents a unified vision transformer for robust multi-agent multi-scale perception. Despite their targeted designs, these methods all follow ”train-from-scratch” paradigm and thus exhibit unstable performance when faced with V2X collaboration challenges, particularly due to sensor data heterogeneity issues [31, 32]. In this paper, we propose a generalizable pretraining framework to enhance the 3D encoders of these backbones. Our framework provides a more robust and versatile representation in the BEV space, directly facilitating the 3D detection task in cooperative perception scenarios from a BEV perspective.
@@ II-B Lidar-based Self-supervised Learning
Representation learning in autonomous driving scenarios [33, 34, 35, 36, 18, 37, 20, 38, 39] has been prevailingly investigated in single-vehicle systems. Stemming from the recent advancement of image reconstruction pretraining methods[40, 41], point cloud pretraining reconstruction methods[42, 43, 44, 45, 46, 47, 19, 18, 37, 20] are also proven effective in improving backbone model’s robustness and generalizability. Recently, this approach has been applied to 3D representation learning of outdoor point clouds. Occupancy-MAE[37] applies a voxel-wise masking strategy to reconstruct masked voxels and predict occupancy. GD-MAE[19] proposes a multi-level transformer architecture and a multi-scale masking strategy with a lightweight generative decoder to recover masked patches. GeoMAE[20] formulates the pretraining target to be geometric feature predictions, such as pyramid centroid, occupancy, surface normal, and curvature of point clouds. BEV-MAE[18] uses a BEV-guided masking strategy to learn BEV feature representations. Notably, while such pretraining methods have shown great potential in developing general feature representations, their improvement is limited due to the restricted perception field caused by perception range and occlusion in single-vehicle systems. Additionally, there is limited discussion on pretraining in multi-agent systems, with the most similar work being CORE[48], which proposes a BEV map reconstruction as an auxiliary task alongside the original perception target. Our work proposes a multi-agent point cloud reconstruction pretraining framework that synergizes effective cooperative pretraining among vehicle and infrastructure agents in V2X cooperative scenarios.
@@ III Method
III-A Overview
The CooPre pretraining pipeline, as shown in Fig.3, involves an early fusion stage where the cooperative agents will transmit their corresponding LiDAR point clouds and metadata to the ego agent. Then we project the vast amount of unlabeled LiDAR point cloud data into a bird-eye-view (BEV) plane, where we perform random masking on the point clouds in predefined BEV grids at a fixed ratio. After that, the unmasked point clouds will be passed to the 3D encoder to obtain a BEV feature, which will be passed to a light decoder to reconstruct masked point clouds from both ego and cooperative agents. At the finetuning stage, the 3D encoder would be taken to perform the downstream cooperative perception task along with the designated fusion network[21].
III-B V2X BEV-guided Masking Strategy
Intuition. In a BEV perspective, directly applying a random masking strategy will often lead to the fact of masking out most of the ego agent’s near-range point cloud without much attention to the far-range point cloud. This makes the pretraining task hard to learn the representation in far ranges. A BEV-guided masking strategy will mitigate this problem by masking the predefined BEV grid that evenly divides the BEV space. However, from a single-agent perspective, such a masking strategy often results in a large number of BEV grids with a sparse point cloud inside each grid, making the self-supervision hard and inaccurate. Multi-agent collaboration becomes powerful as it provides more point cloud occupation from other collaborative agents in that BEV grid with sparse point clouds. Therefore, we extend the BEV-guided masking strategy[18] to a V2X collaborative scenario.
@@ III-B V2X BEV-guided Masking Strategy
V2X LiDAR Point Clouds and Metadata Sharing. In the V2X cooperative scenario, suppose we have ego agent $A_{ego}$ and a set of $N$ cooperative agents $A_{i}$ for $i\in\{1\ldots N\}$ within the communication range, where each agent could be either Connected Autonomous Vehicle (CAV) or infrastructure. During the pretraining stage, each cooperative agent $A_{i}$ shares LiDAR point clouds and metadata information, such as poses, extrinsics, and agent type, to $A_{ego}$. We assume the transmission of LiDAR point clouds and metadata is well-synchronized. Consequently, after projecting the point clouds of each cooperative agent to the ego agent’s coordinate, the perception point cloud field of the ego agent $A_{ego}$ includes its own LiDAR point clouds $P_{ego}$ as well as the LiDAR point clouds $P_{i}$ from each cooperative agent $A_{i}$. We refer to the collection of projected LiDAR point clouds from cooperative agents as $P_{coop}=\bigcup_{i}P_{i}$.
Masking Strategy. For each ego agent $A_{ego}$, given the perception range and the resolution, we construct a BEV space of $X\times Y\times C$ with a set of BEV grids $g_{i,j}$ with a pre-defined voxel size. After that, for each LiDAR point $p_{k}\in P_{ego}\bigcup P_{coop}$, we project it onto a corresponding BEV grid $g_{i,j}$ based on its x, y coordinates. With the help of projected point clouds from cooperative agents, the number of empty BEV grids largely decreases. We then randomly apply a high masking ratio towards non-empty BEV grids, where the point clouds inside the grid will be masked and replaced by learnable point cloud tokens. These tokens will be used alongside with visible point clouds to perform reconstruction target with a light-weight decoder.
III-C Reconstruction Objective
@@
As the number of point clouds varies in each masked BEV grid, we designate the lightweight decoder to output a fixed number of point clouds for reconstruction and utilize the Chamfer distance loss as the learning objective. Specifically, suppose in the masked BEV grid $g_{i,j}$, the reconstruction loss is defined as the following equation between predicted point clouds $\hat{P}$ and groundtruth point clouds $P$:
@@
Note that the groundtruth point clouds in each masked grid contain a mixture of points from $P_{ego}$ and $P_{coop}$.
TABLE I: Performance on V2X-Real dataset. Green denotes the improvements over the baseline AttFuse backbone.
@@
CooPre 71.5 +3.8 70.2 +4.4 46.9 +1.9 28.0 +0.4 61.9 +6.4 58.3 +6.3 60.1 +4.0 52.2 +3.7
@@
CooPre 71.4 +4.5 70.1 +4.8 49.7 +3.7 33.7 +3.3 57.5 +3.7 55.5 +5.3 59.5 +4.0 53.1 +4.5
@@
CooPre 86.2 +1.7 84.1 +2.0 60.5 -0.5 39.3 -1.3 61.5 +1.6 61.0 +1.9 69.4 +0.9 61.4 +0.8
@@
CooPre 84.6 +1.2 82.7 +0.8 66.5 -0.5 45.2 +0.4 65.2 +1.6 62.5 -0.3 72.1 +0.8 63.5 +0.3
Figure 4: Comparison of perception performance with different training paradigms in V2X-Real VC and V2V datasets. In comparison to train-from-scratch and single-agent pretraining methods, our multi-agent pretraining yields more robust perception capabilities in scenarios involving occlusions and long-range perception. Particularly, the occurrence of false perceptions significantly decreases in far-range scenarios where point cloud features are sparse. Blue and red 3D bounding boxes correspond to the ground-truth and detection outputs, respectively.
TABLE II: Performance on V2V4Real dataset. Green denotes the improvements over the baseline AttFuse backbone.
@@
CooPre 74.3/49.3 +3.1/5.1 87.4/60.2 +0.1/4.7 63.4/45.2 +7.8/5.4 55.6/32.0 +4.5/6.3 66.1/39.9 +2.3/5.0 85.9/57.0 +0.2/5.6 54.3/34.9 +6.5/3.9 36.1/14.0 +2.7/3.5
TABLE III: Performance on OPV2V dataset. Green denotes the improvements over the baseline AttFuse backbone.
@@
CooPre 91.7 +2.4 86.6 +4.0 88.7 +1.2 80.0 +4.0

### v2

@@ Abstract
Existing Vehicle-to-Everything (V2X) cooperative perception methods rely on accurate multi-agent 3D annotations. Nevertheless, it is time-consuming and expensive to collect and annotate real-world data, especially for V2X systems. In this paper, we present a self-supervised learning framwork for V2X cooperative perception, which utilizes the vast amount of unlabeled 3D V2X data to enhance the perception performance. Specifically, multi-agent sensing information is aggregated to form a holistic view and a novel proxy task is formulated to reconstruct the LiDAR point clouds across multiple connected agents to better reason multi-agent spatial correlations. Besides, we develop a V2X bird-eye-view (BEV) guided masking strategy which effectively allows the model to pay attention to 3D features across heterogeneous V2X agents (i.e., vehicles and infrastructure) in the BEV space. Noticeably, such a masking strategy effectively pretrains the 3D encoder with a multi-agent LiDAR point cloud reconstruction objective and is compatible with mainstream cooperative perception backbones. Our approach, validated through extensive experiments on representative datasets (i.e., V2X-Real, V2V4Real, and OPV2V) and multiple state-of-the-art cooperative perception methods (i.e., AttFuse, F-Cooper, and V2X-ViT), leads to a performance boost across all V2X settings. Notably, CooPre achieves a 4% mAP improvement on V2X-Real dataset and surpasses baseline performance using only 50% of the training data, highlighting its data efficiency. Additionally, we demonstrate the framework’s powerful performance in cross-domain transferability and robustness under challenging scenarios. The code will be made publicly available at https://github.com/ucla-mobility/CooPre.
@@ I Introduction
Figure 1: Illustration of our CooPre framework. Given multi-agent LiDAR input, CooPre pretrains the 3D encoder via LiDAR point cloud reconstruction. During the finetuning stage, the initialized 3D encoder will be used for 3D detection.
To this end, we propose an effective self-supervised multi-agent Cooperative Pretraining framework (CooPre) that enables the model to learn meaningful prior representation of the holistic 3D environment before the perception task, as illustrated in Fig. 1. Our framework leverages the benefits of unlabeled LiDAR point cloud data transmitted from different agents, allowing the model to reconstruct the point cloud location and learn essential prior knowledge of scenarios (e.g., intersections or corridors) and LiDAR sensor distributions (e.g., range, placement position, and sparsity) of each agent from a bird-eye-view (BEV) perspective. In the enlarged perception field, this heterogeneous reconstruction task also helps mitigate issues related to sparse feature points in far-range and occlusion scenarios. In addition, this method can be seamlessly integrated with state-of-the-art cooperative perception methods to improve perception capability. Compared to the other two training paradigms in Fig. 2, our pretraining framework is annotation-free, which is the main contribution to representation learning in V2X cooperative perception.
Through extensive experiments and ablation studies across three representative V2X datasets (i.e., V2X-Real [1], V2V4Real [8], and OPV2V [18]) and three different V2X fusion methods [18, 19, 2], we demonstrate the effectiveness of CooPre in V2X cooperative perception. Our findings show that: 1) CooPre boosts perception capabilities in scenarios involving occlusions and long-range perception, outperforming train-from-scratch and single-agent pretraining methods [20, 21]; 2) CooPre enables the model to learn important geometric and topological representations under heterogeneous sensor configurations, allowing better recognition of rigid-body objects like cars and trucks; 3) the learned representations from CooPre enhance data efficiency and demonstrate strong generalizability to unseen distributions, making them particularly advantageous for domain adaptation, especially in data scarcity scenarios. Moreover, our study underscores the critical role of well-learned 3D representations as a promising complement to task-specific design optimizations in V2X cooperative perception.
@@ II-A Cooperative Perception
Single-vehicle systems struggle with occlusions and long-distance perception in complex traffic environments due to the limitation of the perception range of LiDAR devices. Cooperative systems, on the other hand, enhance detection performance by sharing raw data (Early Fusion), detection outputs (Late Fusion), or intermediate bird-eye-view (BEV) representations (Intermediate Fusion) among connected agents [18, 22, 23]. With the recent advancement of a variety of simulation and real-world dataset curations [18, 8, 3, 24, 4, 1, 5], many literature [18, 2, 25, 22, 19, 23, 26, 27] have been discussing the algorithmic designs of collaborative modes from vehicle-to-vehicle (V2V) collaboration to vehicle-to-everything (V2X) collaboration in different traffic scenarios. Noticeably, the intermediate fusion strategy has been the primary direction since it achieves the best trade-off between accuracy and bandwidth requirements. After applying a 3D encoder [28, 29, 30] to the input LiDAR feature, the intermediate fusion strategy involves projecting 3D features to BEV features where agents will perform interactions before final detection results. For example, AttFuse [18] utilizes a simple agent-wise single-head attention to fuse all features, whereas V2X-ViT [2] presents a unified vision transformer for robust multi-agent multi-scale perception. Despite their targeted designs, these methods all follow ”train-from-scratch” paradigm and thus exhibit unstable performance when faced with V2X collaboration challenges, particularly due to sensor data heterogeneity issues [31, 32]. In this paper, we propose a model-agnostic pretraining framework to enhance the perception capability of these cooperative perception methods.
@@ II-B Lidar-based Self-supervised Learning
Representation learning in autonomous driving [33, 34, 35, 36, 20, 37, 38, 39, 40, 21] has been prevailingly investigated in single-vehicle systems. Stemming from the recent advancement of image reconstruction pretraining methods [41, 42], point cloud pretraining reconstruction methods [43, 44, 45, 46, 47, 48, 49, 20, 37, 38] are also proven effective in improving backbone model’s robustness and generalizability. Recently, this approach has been applied to 3D representation learning of outdoor point clouds. Occupancy-MAE [37] applies a voxel-wise masking strategy to reconstruct masked voxels and predict occupancy. GD-MAE [49] proposes a multi-level transformer architecture and a multi-scale masking strategy with a lightweight generative decoder to recover masked patches. GeoMAE [38] formulates the pretraining target to be geometric feature predictions, such as pyramid centroid, occupancy, surface normal, and curvature of point clouds. BEV-MAE [20] uses a BEV-guided masking strategy to learn BEV feature representations. Notably, while such pretraining methods have shown great potential in developing general feature representations, their improvement is limited due to the restricted perception field caused by perception range and occlusion in single-vehicle systems. Additionally, there is limited discussion on pretraining in multi-agent systems, with the most similar work being CORE [50], which proposes a BEV map reconstruction as an auxiliary task alongside the original perception target. Our work, on the other hand, provides the well-learned BEV representations that facilitate perception performance without additional computation efforts during the finetuning stage.
@@ III Method
III-A Framework Overview
Our CooPre framework, as illustrated in Fig. 3, is designed to facilitate 3D representation learning across heterogeneous V2X agents. To ensure robust cross-agent feature learning and mitigate feature sparsity issues, we propose three key design elements. First, we leverage early fusion to integrate LiDAR point cloud from collaborative agents, effectively expanding the ego agent’s perception field. This design mitigates the sparse feature issue and enables the subsequent LiDAR point cloud reconstruction objective to be applied to both the ego and collaborative agents. Second, instead of reconstructing the point cloud in voxel space, we reconstruct masked information on the BEV plane. This design enhances the model’s feature learning process from a BEV perspective. Third, we require the model to reconstruct masked LiDAR point clouds from both the ego and collaborative agents, allowing the 3D encoder to learn the multi-agent LiDAR sensor distribution regardless of the ego agent’s perspective. Notably, our pretraining operates in BEV representation space, making it compatible with mainstream cooperative fusion methods [18, 19, 2] and widely used 3D encoder backbones [30, 28] during finetuning.
III-B Multi-agent Reconstruction Pretraining
@@
Early Fusion of Multi-agent LiDAR point cloud. We formulate the process of early fusion of raw LiDAR point cloud data across V2X agents as follows. Suppose we have ego agent $A_{ego}$ and a set of $N$ cooperative agents $A_{i}$ for $i\in\{1\ldots N\}$ within the communication range, where each agent could be either Connected Autonomous Vehicle (CAV) or infrastructure. During the pretraining stage, each cooperative agent $A_{i}$ shares LiDAR point clouds and metadata information, such as poses, extrinsics, and agent type, to $A_{ego}$. We assume the transmission of LiDAR point clouds and metadata is well-synchronized. Consequently, after projecting the point clouds of each cooperative agent to the ego agent’s coordinate, the perception point cloud field of the ego agent $A_{ego}$ includes its own LiDAR point clouds $P_{ego}$ as well as the LiDAR point clouds $P_{i}$ from each cooperative agent $A_{i}$. We refer to the collection of projected LiDAR point clouds from cooperative agents as $P_{coop}=\bigcup_{i}P_{i}$. In order for the encoder to learn invariant features from raw point cloud data from each agent, we perform data augmentations, including scaling, rotation, flip and downsampling [28]. These augmentations enhance the model’s generalization capabilities during pretraining, allowing it to effectively accomplish the pretraining objectives under varying conditions and viewpoints. Thus, the collection of original and augmented multi-agent LiDAR point cloud becomes our pretraining data corpus.
V2X BEV-guided Masking Strategy. We design a V2X BEV-guided masking strategy to enhance feature learning from a BEV perspective. For each ego agent $A_{ego}$, we suppose the shape of obtained BEV features from raw LiDAR input is $X\times Y\times C$. According to this shape, we define a BEV plane of size $X\times Y$ to divide up the plane and obtain BEV grids. We ensure that the size of masked BEV grids matches the resolution of BEV features obtained from 3D encoder. Each BEV grid $g_{i,j}$ would be used to determine whether the points inside the grid would be masked or not. After that, for each LiDAR point $p_{k}\in P_{ego}\bigcup P_{coop}$, we project it onto a corresponding BEV grid $g_{i,j}$ based on $p_{k}$’s x, y coordinates. With the help of projected point clouds from cooperative agents, the number of empty BEV grids largely decreases. We then randomly apply a high masking ratio towards non-empty BEV grids, where the point clouds inside the grid will be masked. Note that for a single-agent approach [20], this strategy often leads to numerous BEV grids with sparse point clouds, making self-supervision both difficult and imprecise. In contrast, multi-agent collaboration proves advantageous, as it enriches the sparsely populated BEV grids with point cloud data from other collaborative agents.
Decoder design. The decoder design is flexible and independent from downstream task since only the 3D encoder would be taken for finetuning. Unlike [49], we utilize a lightweight decoder with one convolution layer for training efficiency. The decoder would take unmasked point cloud as input and output a fixed number of LiDAR point clouds for the multi-agent reconstruction objective.
@@
Multi-agent LiDAR point cloud Reconstruction Objective. As the ground-truth number of point clouds varies in each masked BEV grid, we utilize the Chamfer distance loss as the learning objective to measure the difference between predicted point clouds $\hat{P}$ and ground-truth point clouds $P$. Specifically, suppose in the masked BEV grid $g_{i,j}$, the reconstruction loss is defined as the following equation:
@@
We enforce the model to reconstruct the LiDAR point cloud from both ego and cooperative agent $P_{ego}$ and $P_{coop}$. We will show in the ablation study that such design is crucial for far-range feature learning across heterogeneous agents.
TABLE I: Performance on V2X-Real dataset.
@@
F-Cooper [19] 64.9 62.7 45.5 26.7 39.8 35.1 50.1 41.5
V2X-ViT [2] 63.3 61.2 42.2 24.8 38.4 36.0 48.0 40.0
CooPre (AttFuse) 71.5 +3.8 70.2 +4.4 46.9 +1.9 28.0 +0.4 61.9 +6.4 58.3 +6.3 60.1 +4.0 52.2 +3.7
CooPre (F-Cooper) 67.0 +2.1 65.7 +3.0 43.6 -1.9 25.4 -1.3 49.3 +9.5 44.9 +9.8 53.3 +3.2 45.3 +3.8
CooPre (V2X-ViT) 66.1 +2.8 64.5 +3.3 45.3 +3.1 27.7 +2.9 43.5 +5.1 40.1 +4.1 51.7 +3.7 44.1 +4.1
@@
F-Cooper [19] 64.6 63.0 41.9 26.9 48.9 44.8 51.8 44.9
V2X-ViT [2] 62.6 58.7 47.5 30.5 49.3 47.0 53.1 45.4
CooPre (AttFuse) 71.4 +4.5 70.1 +4.8 49.7 +3.7 33.7 +3.3 57.5 +3.7 55.5 +5.3 59.5 +4.0 53.1 +4.5
CooPre (F-Cooper) 68.6 +4.0 67.0 +4.0 42.0 +0.1 26.4 -0.5 51.0 +2.1 48.1 +3.3 53.9 +2.1 47.2 +2.3
CooPre (V2X-ViT) 66.9 +4.3 64.5 +5.8 46.7 -0.8 29.2 -1.3 54.1 +4.8 52.3 +5.3 55.9 +2.8 48.7 +3.3
@@
F-Cooper [19] 79.1 76.2 57.0 35.8 52.0 45.7 62.7 52.6
V2X-ViT [2] 81.0 76.9 44.9 29.2 37.3 36.8 54.4 47.6
CooPre (AttFuse) 86.2 +1.7 84.1 +2.0 60.5 -0.5 39.3 -1.3 61.5 +1.6 61.0 +1.9 69.4 +0.9 61.4 +0.8
CooPre (F-Cooper) 84.1 +5.0 80.7 +4.5 58.4 +1.4 33.9 -1.9 51.8 -0.2 50.7 +5.0 64.8 +2.1 55.1 +2.5
CooPre (V2X-ViT) 81.4 +0.4 77.3 +0.4 45.8 +0.9 29.1 -0.1 37.4 +0.1 37.2 +0.4 54.8 +0.4 47.9 +0.3
@@
F-Cooper [19] 85.1 82.0 66.1 44.3 51.3 50.8 67.5 59.0
V2X-ViT [2] 82.0 79.5 68.9 45.4 59.0 57.0 69.9 60.6
CooPre (AttFuse) 84.6 +1.2 82.7 +0.8 66.5 -0.5 45.2 +0.4 65.2 +1.6 62.5 -0.3 72.1 +0.8 63.5 +0.3
CooPre (F-Cooper) 85.4 +0.3 82.1 +0.1 66.6 +0.5 45.1 +0.8 51.9 +0.6 51.1 +0.3 68.0 +0.5 59.4 +0.4
CooPre (V2X-ViT) 84.0 +2.0 82.4 +2.9 66.9 -2.0 44.4 -1.0 59.1 +0.1 58.7 +1.7 70.0 +0.1 61.9 +1.3
Figure 4: Illustration of enhanced detection performance in long-range and occlusion scenarios. Our multi-agent pretraining framework yields more robust perception capabilities in scenarios involving occlusions and long-range perception. Particularly, the occurrence of false perceptions significantly decreases in far-range scenarios where point cloud features are sparse. Blue and red 3D bounding boxes correspond to the ground-truth and detection outputs, respectively. Orange boxes denote the zoomed-in view of detection results.
TABLE II: Performance on V2V4Real dataset.
@@
F-Cooper [19] 70.6/41.8 84.4/52.8 59.9/37.8 51.5/26.2 62.2/34.1 82.5/50.0 51.7/30.9 31.9/10.7
V2X-ViT [2] 71.7/43.3 87.8/59.5 58.6/34.1 50.4/23.0 60.0/32.9 83.3/51.3 47.4/25.7 26.3/9.0
CooPre (AttFuse) 74.3/49.3 +3.1/+5.1 87.4/60.2 +0.1/+4.7 63.4/45.2 +7.8/+5.4 55.6/32.0 +4.5/+6.3 66.1/39.9 +2.3/+5.0 85.9/57.0 +0.2/+5.6 54.3/34.9 +6.5/+3.9 36.1/14.0 +2.7/+3.5
CooPre (F-Cooper) 71.4/43.4 +0.8/+1.6 86.1/53.2 +1.7/+0.4 63.0/39.6 +3.1/+1.8 48.2/29.5 -3.3/+3.3 63.7/34.8 +1.5/+0.7 84.7/50.3 +2.2/+0.3 55.0/31.1 +3.3/+0.2 29.7/11.9 -2.2/+0.8
CooPre (V2X-ViT) 73.1/43.9 +1.4/+0.6 88.9/58.1 +1.1/-1.4 61.6/36.7 +3.0/+2.6 50.7/26.0 +0.3/+3.0 63.8/38.2 +3.8/+5.3 86.7/58.3 +3.4/+7.0 51.1/30.7 +3.7/+5.0 29.8/11.3 +3.5/+2.3
TABLE III: Performance on OPV2V dataset.
@@
F-Cooper [19] 87.8 81.7 89.0 79.7
V2X-ViT [2] 89.8 83.7 88.7 80.2
CooPre (AttFuse) 91.7 +2.4 86.6 +4.0 88.7 +1.2 80.0 +4.0
CooPre (F-Cooper) 89.0 +1.2 83.4 +1.7 90.2 +1.2 82.0 +2.3
CooPre (V2X-ViT) 91.9 +2.1 86.5 +2.8 89.9 +1.2 82.4 +2.2